OpenStudy (swaqqout_kid1):

Seeley is using a cart to move several large boxes at a time. If each box weighs 17 lb and the cart weighs 12 lb, how much total weight must he move if he loads 4 boxes onto the cart? @sangya21 @sammixboo

OpenStudy (anonymous):

total weight = 17(x) + 12 for x = 4 w = 17(4) + 12 we did this kind of questions http://openstudy.com/users/swaqqout_kid1#/updates/547b84aae4b082a9810f2d44

sammixboo (sammixboo):

Yeppers! we have \(\rm (17b)~+~12\) b = number of boxes = 4 So plug in the value\ \(\rm (17\times 4)~+~12\) Can you tell me what is \(\rm 17\times4?\)

OpenStudy (swaqqout_kid1):

68

sammixboo (sammixboo):

Correct! \(\rm (17\times 4)~+~12\) \(\rm 68~+~12\) I bet you know what to do next :P

OpenStudy (swaqqout_kid1):

80

sammixboo (sammixboo):

Correct! So the cart all together will weigh 80 lbs

OpenStudy (swaqqout_kid1):

hea u message me